OpenStudy (unklerhaukus):

I would appreciate some assistance \[\frac{\text d^2 y}{\text dx^2}+n^2y=0\]

OpenStudy (unklerhaukus):

i can solve easily using second order techniques: \[\frac{\text d^2 y}{\text dx^2}+n^2y=0\]\[m^2+n^2=0\]\[m=\pm in\]\[y=A\cos nx+B\sin nx\]

OpenStudy (unklerhaukus):

but im finding it hard with first order techniques: \[\frac{\text d^2 y}{\text dx^2}+n^2y=0\]\[\text {let } \frac{\text dy}{\text dx}=p\]\[\frac{\text d^2y}{\text dx^2}=\frac{\text dp}{\text dx}=\frac{\text dp}{\text dy}\frac{\text dy}{\text dx}=\frac{\text dp}{\text dy}p\]\[\frac{\text dp}{\text dy}+n^2y=0\]\[\frac{\text dp}{\text dy}=-n^2y\]\[\int \text dp=-n^2\int y\text dy\]\[p=-n^2\frac{y^2}{2}+c\]\[\frac{\text dy}{\text dx}=-n^2\frac{y^2}{2}+c\]\[-2\frac{\text dy}{n^2y^2+c}=\text dx\]\[-2\int\frac{\text dy}{n^2y^2+c}=\int\text dx\]\[-2\left(\frac{\arctan{\frac{ny}{\sqrt c}}}{\sqrt cn}\right)=x+d\]\[\arctan{\frac{ny}{\sqrt c}}=\frac{\sqrt c n(x+d)}2\]\[\frac{ny}{\sqrt c}=\tan\left(\frac{\sqrt c n(x+d)}2\right)\]\[\frac{ny}{\sqrt c}=\frac{\sin\left(\frac{\sqrt c n(x+d)}2\right)}{\cos \left(\frac{\sqrt c n(x+d)}2\right)}\]

OpenStudy (anonymous):

I heard 42 was the answer to everything O.O

OpenStudy (unklerhaukus):

im not looking of an answer im looking for a method

OpenStudy (anonymous):

lol have you tried PEMDAS ;D

OpenStudy (anonymous):

better yet; try elmo counts

OpenStudy (unklerhaukus):

im not evaluating @rebeccaskell94

OpenStudy (unklerhaukus):

i cant count any elmos here,

OpenStudy (unklerhaukus):

i think i can see a mistake in the working